OpenStudy (anonymous):

What is the refractive index of a medium when the light strikes the medium after traveling through air at an angle of 23.3° and the refracted angle is 14.6°? A. 1.33 B. 1.36 C. 1.44 D. 1.57 E. 1.91

OpenStudy (michele_laino):

 please apply the law of snell, namely: \[n _{air}*\sin \theta _{1}=n _{medium}\sin \theta _{2}\]

OpenStudy (anonymous):

Law of who now? also i don't have a calc.

OpenStudy (michele_laino):

law of Snell

OpenStudy (michele_laino):

please use Windows calc!

OpenStudy (michele_laino):

please note that your refractive index is equals to: \[\frac{ m _{medium} }{ n _{air} }\] referring to the formula that I wrote above

OpenStudy (michele_laino):

oops sorry: \[\frac{ n _{medium} }{ n _{air} }\]

OpenStudy (anonymous):

I don't understand what doesn n stand for and what is the medium?

OpenStudy (anonymous):

@HarshTheAwesome

OpenStudy (michele_laino):

your refractive index is a relative refractive index respect with to air, it is equals to: \[\frac{ n _{medium} }{ n _{air} }=\frac{ \sin \theta _{1} }{ \sin \theta _{2} }\]

OpenStudy (anonymous):

The answer choices are weird

OpenStudy (michele_laino):

with: \[\theta _{1}=23.3,\theta _{2}=14.6\]

OpenStudy (anonymous):

My bad, it's 1.57

OpenStudy (anonymous):

sin(23.3)/sin(14.6)

OpenStudy (michele_laino):

@Odette32691148 please try to substitute your numerica data in the formula above!

OpenStudy (michele_laino):

try please! \[\frac{ n _{medium} }{ n _{air} }=\frac{ \sin(23.3) }{ \sin(14.6) }=...\]

OpenStudy (michele_laino):

use Windows calc, please!

OpenStudy (michele_laino):

select "scientific" view from menu, please!